A plane leaves an airport traveling at 400 mph in the direction N 45° E. A wind is blowing at 40 mph in the direction N 45° W. What is the ground speed of the plane?
1
Expert's answer
2020-12-03T19:09:27-0500

The figure below represents a right angled triangle. Using pythagoras theorem we have that

"GS^2 = 400^2 + 40^2" where GS represents the ground speed, this implies that

"GS = \sqrt{400^2 + 40^2}" , therefore

GS = 401.995mph

Using sine rule

"\frac{A}{sinA} = \frac{B}{sinB}"


= "\frac{401.995}{sin90} = \frac{40}{sinx}" ,therefore "x=5.9105^0"

this implies the direction of the ground speed is "50.9105^0"

Therefore the ground speed is 401.995mph at "N50.91^0E"